OpenStudy (anonymous):

Use Subsitution: 7x+2y=-6 -14x-4y=-2

myininaya (myininaya):

how about multiplying the first equation by 2. and solving it for 4y and plugging that 4y into the equation equation to solve for x so we have 14x+4y=-12 => 4y=-12-14x -14x-4y=-2 => -14x-(-12-14x)=-2 => 12 =-2 (not true) also if we write both of them in this form y=mx+b y=-14x/4-12/4 -4y=14x-2 => y=-14x/4+2/4 same slope but different y intercept => lines are parallel => no solution

OpenStudy (anonymous):

thank you so much!